# Proactive and Reactive Constraint Programming for Stochastic Project Scheduling with Maximal Time-Lags

### v1

@@ Appendix D Reproducibility
[288] -8.783 (*) [285] -10.58 (*) [284] -13.068 (*) [122] -4.786 (*) [146] -5.008 (*) [141] -1.441 (0.15)
[288] 0.576 (*) [285] 0.653 (*) [284] 0.82 (*) [23] 1.0 (*) [146] 0.836 (*) [141] 0.702 (*)
Table 18: Pairwise comparison on time offline for noise factor c=2. Using a Wilcoxon test and a proportion test. Including all instances for which at least one of the two methods found a feasible solution. Note that the ordering matters: the first method showed is the better of the two in each pair method 1 - method 2 according to Wilcoxon. Each cell shows on the first row [nr pairs] z-value (p-value) of the Wilcoxon test with (*) for p $<$ 0.05. Each cell shows on the second row [nr pairs] ratio of wins (p-value) with (*) for p $<$ 0.05
[233] nan (nan) [205] -229.689 (*) [230] -59.39 (*) [206] -230.812 (*) [231] -59.404 (*) [204] -17.066 (*)
$\mathsf{pro}_{0.9}$: 1.0 $\mathsf{react}$: 0.15 $\mathsf{react}$: 0.35 $\mathsf{pro}_{0.9}$: 0.15 $\mathsf{pro}_{0.9}$: 0.35 $\mathsf{pro}_{\mathsf{SAA}}$: 0.61
$\mathsf{react}$: 1.0 $\mathsf{stnu}$: 1.85 $\mathsf{pro}_{\mathsf{SAA}}$: 1.65 $\mathsf{stnu}$: 1.85 $\mathsf{pro}_{\mathsf{SAA}}$: 1.65 $\mathsf{stnu}$: 1.39
[176] nan (nan) [170] -82.607 (*) [174] -63.732 (*) [170] -82.607 (*) [173] -63.382 (*) [172] 0.379 (0.705)
$\mathsf{pro}_{0.9}$: 1.0 $\mathsf{react}$: 0.21 $\mathsf{react}$: 0.23 $\mathsf{pro}_{0.9}$: 0.21 $\mathsf{pro}_{0.9}$: 0.23 $\mathsf{stnu}$: 1.01
$\mathsf{react}$: 1.0 $\mathsf{stnu}$: 1.79 $\mathsf{pro}_{\mathsf{SAA}}$: 1.77 $\mathsf{stnu}$: 1.79 $\mathsf{pro}_{\mathsf{SAA}}$: 1.77 $\mathsf{pro}_{\mathsf{SAA}}$: 0.99
[160] nan (nan) [140] -33.765 (*) [152] -40.336 (*) [138] -33.253 (*) [150] -39.904 (*) [148] -2.168 (*)
$\mathsf{pro}_{0.9}$: 1.0 $\mathsf{react}$: 0.34 $\mathsf{react}$: 0.33 $\mathsf{pro}_{0.9}$: 0.34 $\mathsf{pro}_{0.9}$: 0.33 $\mathsf{stnu}$: 0.9
$\mathsf{react}$: 1.0 $\mathsf{stnu}$: 1.66 $\mathsf{pro}_{\mathsf{SAA}}$: 1.67 $\mathsf{stnu}$: 1.66 $\mathsf{pro}_{\mathsf{SAA}}$: 1.67 $\mathsf{pro}_{\mathsf{SAA}}$: 1.1
[148] 26.653 (*) [141] 27.168 (*) [146] -0.952 (0.343) [155] nan (nan) [150] -45.457 (*) [145] -44.29 (*)
$\mathsf{stnu}$: 1.7 $\mathsf{stnu}$: 1.72 $\mathsf{stnu}$: 0.95 $\mathsf{pro}_{0.9}$: 1.0 $\mathsf{pro}_{0.9}$: 0.24 $\mathsf{react}$: 0.24
$\mathsf{pro}_{0.9}$: 0.3 $\mathsf{react}$: 0.28 $\mathsf{pro}_{\mathsf{SAA}}$: 1.05 $\mathsf{react}$: 1.0 $\mathsf{pro}_{\mathsf{SAA}}$: 1.76 $\mathsf{pro}_{\mathsf{SAA}}$: 1.76
[114] 14.765 (*) [94] 15.183 (*) [114] -3.125 (*) [99] nan (nan) [94] -42.151 (*) [76] -42.944 (*)
$\mathsf{stnu}$: 1.55 $\mathsf{stnu}$: 1.61 $\mathsf{stnu}$: 0.83 $\mathsf{pro}_{0.9}$: 1.0 $\mathsf{pro}_{0.9}$: 0.21 $\mathsf{react}$: 0.18
$\mathsf{pro}_{0.9}$: 0.45 $\mathsf{react}$: 0.39 $\mathsf{pro}_{\mathsf{SAA}}$: 1.17 $\mathsf{react}$: 1.0 $\mathsf{pro}_{\mathsf{SAA}}$: 1.79 $\mathsf{pro}_{\mathsf{SAA}}$: 1.82
Table 19: Magnitude test on time offline for noise factor c=2. Using a pairwise t-test, including all instances for which both methods found a feasible solution. Each cell shows on the first row [nr pairs] t-stat (p-value) with (*) for p $<$ 0.05 and on the second row the normalized average of method 1 and on the third row the normalized average of method 2.
@@
[400] -0.719 (0.472) [400] -17.331 (*) [400] -17.331 (*) [400] -17.331 (*) [400] -17.331 (*) [400] -12.488 (*)
[391] 0.453 (0.069) [400] 1.0 (*) [400] 1.0 (*) [400] 1.0 (*) [400] 1.0 (*) [400] 0.925 (*)
Table 20: Pairwise comparison on time online for noise factor c=1. Using a Wilcoxon test and a proportion test. Including all instances for which at least one of the two methods found a feasible solution. Note that the ordering matters: the first method showed is the better of the two in each pair method 1 - method 2. Each cell shows on the first row [nr pairs] z-value (p-value) of the Wilcoxon test with (*) for p $<$ 0.05. Each cell shows on the second row [nr pairs] proportion (p-value) with (*) for p $<$ 0.05
@@
[340] 0.679 (0.498) [260] -109.014 (*) [260] -106.917 (*) [340] -4393.475 (*) [340] -4277.995 (*) [260] -1151.565 (*)
$\mathsf{pro}_{0.9}$: 1.02 $\mathsf{pro}_{0.9}$: 0.05 $\mathsf{pro}_{\mathsf{SAA}}$: 0.05 $\mathsf{pro}_{0.9}$: 0.0 $\mathsf{pro}_{\mathsf{SAA}}$: 0.0 $\mathsf{stnu}$: 0.04
$\mathsf{pro}_{\mathsf{SAA}}$: 0.98 $\mathsf{stnu}$: 1.95 $\mathsf{stnu}$: 1.95 $\mathsf{react}$: 2.0 $\mathsf{react}$: 2.0 $\mathsf{react}$: 1.96
@@
[270] -0.663 (0.508) [230] -285.563 (*) [230] -282.464 (*) [270] -7311.952 (*) [270] -7470.984 (*) [230] -605.38 (*)
$\mathsf{pro}_{0.9}$: 0.97 $\mathsf{pro}_{0.9}$: 0.03 $\mathsf{pro}_{\mathsf{SAA}}$: 0.03 $\mathsf{pro}_{0.9}$: 0.0 $\mathsf{pro}_{\mathsf{SAA}}$: 0.0 $\mathsf{stnu}$: 0.07
$\mathsf{pro}_{\mathsf{SAA}}$: 1.03 $\mathsf{stnu}$: 1.97 $\mathsf{stnu}$: 1.97 $\mathsf{react}$: 2.0 $\mathsf{react}$: 2.0 $\mathsf{react}$: 1.93
[320] -0.156 (0.876) [280] -789.246 (*) [280] -786.726 (*) [320] -14638.229 (*) [320] -14425.009 (*) [280] -338.664 (*)
$\mathsf{pro}_{0.9}$: 0.99 $\mathsf{pro}_{0.9}$: 0.02 $\mathsf{pro}_{\mathsf{SAA}}$: 0.02 $\mathsf{pro}_{0.9}$: 0.0 $\mathsf{pro}_{\mathsf{SAA}}$: 0.0 $\mathsf{stnu}$: 0.09
$\mathsf{pro}_{\mathsf{SAA}}$: 1.01 $\mathsf{stnu}$: 1.98 $\mathsf{stnu}$: 1.98 $\mathsf{react}$: 2.0 $\mathsf{react}$: 2.0 $\mathsf{react}$: 1.91
[370] -0.795 (0.427) [330] -3951.584 (*) [330] -3976.441 (*) [370] -30723.506 (*) [370] -31447.766 (*) [330] -142.758 (*)
$\mathsf{pro}_{\mathsf{SAA}}$: 0.97 $\mathsf{pro}_{0.9}$: 0.01 $\mathsf{pro}_{\mathsf{SAA}}$: 0.01 $\mathsf{pro}_{0.9}$: 0.0 $\mathsf{pro}_{\mathsf{SAA}}$: 0.0 $\mathsf{stnu}$: 0.15
$\mathsf{pro}_{0.9}$: 1.03 $\mathsf{stnu}$: 1.99 $\mathsf{stnu}$: 1.99 $\mathsf{react}$: 2.0 $\mathsf{react}$: 2.0 $\mathsf{react}$: 1.85
[400] -0.065 (0.948) [370] -56517.571 (*) [370] -48810.689 (*) [388] -83435.86 (*) [388] -78013.468 (*) [358] -60.202 (*)
$\mathsf{pro}_{\mathsf{SAA}}$: 1.0 $\mathsf{pro}_{0.9}$: 0.0 $\mathsf{pro}_{\mathsf{SAA}}$: 0.0 $\mathsf{pro}_{0.9}$: 0.0 $\mathsf{pro}_{\mathsf{SAA}}$: 0.0 $\mathsf{stnu}$: 0.23
$\mathsf{pro}_{0.9}$: 1.0 $\mathsf{stnu}$: 2.0 $\mathsf{stnu}$: 2.0 $\mathsf{react}$: 2.0 $\mathsf{react}$: 2.0 $\mathsf{react}$: 1.77
Table 21: Magnitude t-test on time online for noise factor c=1 (double hits). Each cell shows on the first row [nr pairs] t-stat (p-value) with (*) for p $<$ 0.05 and on the second row the normalized average of method 1 and on the third row the normalized average of method 2.
[235] -0.384 (0.701) [258] -9.554 (*) [258] -9.212 (*) [234] -13.262 (*) [235] -12.623 (*) [258] -8.801 (*)
[66] 0.47 (0.712) [258] 0.907 (*) [258] 0.899 (*) [234] 1.0 (*) [235] 0.987 (*) [258] 0.891 (*)
[182] -1.204 (0.229) [216] -4.456 (*) [217] -4.913 (*) [177] -11.279 (*) [182] -10.957 (*) [217] -11.457 (*)
[79] 0.43 (0.261) [216] 0.815 (*) [217] 0.825 (*) [177] 0.994 (*) [182] 0.984 (*) [217] 0.968 (*)
@@
[146] -0.482 (0.63) [288] -8.467 (*) [284] -9.06 (*) [122] -9.585 (*) [141] -6.136 (*) [285] -14.285 (*)
[144] 0.458 (0.359) [288] 0.576 (*) [284] 0.585 (*) [122] 1.0 (*) [141] 0.837 (*) [285] 0.982 (*)
Table 22: Pairwise comparison on time online for noise factor c=2. Using a Wilcoxon test and a proportion test. Including all instances for which at least one of the two methods found a feasible solution. Note that the ordering matters: the first method showed is the better of the two in each paier method 1 - method 2. Each cell shows on the first row [nr pairs] z-value (p-value) of the Wilcoxon test with (*) for p $<$ 0.05. Each cell shows on the second row [nr pairs] proportion (p-value) with (*) for p $<$ 0.05
[231] 0.391 (0.697) [206] -92.546 (*) [204] -91.532 (*) [233] -4426.055 (*) [230] -4653.225 (*) [205] -1399.699 (*)
$\mathsf{pro}_{0.9}$: 1.01 $\mathsf{pro}_{0.9}$: 0.06 $\mathsf{pro}_{\mathsf{SAA}}$: 0.06 $\mathsf{pro}_{0.9}$: 0.0 $\mathsf{pro}_{\mathsf{SAA}}$: 0.0 $\mathsf{stnu}$: 0.03
$\mathsf{pro}_{\mathsf{SAA}}$: 0.99 $\mathsf{stnu}$: 1.94 $\mathsf{stnu}$: 1.94 $\mathsf{react}$: 2.0 $\mathsf{react}$: 2.0 $\mathsf{react}$: 1.97
[173] 0.77 (0.443) [170] -241.897 (*) [172] -262.612 (*) [176] -7787.819 (*) [174] -8817.279 (*) [170] -935.848 (*)
$\mathsf{pro}_{0.9}$: 1.03 $\mathsf{pro}_{0.9}$: 0.03 $\mathsf{pro}_{\mathsf{SAA}}$: 0.03 $\mathsf{pro}_{0.9}$: 0.0 $\mathsf{pro}_{\mathsf{SAA}}$: 0.0 $\mathsf{stnu}$: 0.06
$\mathsf{pro}_{\mathsf{SAA}}$: 0.97 $\mathsf{stnu}$: 1.97 $\mathsf{stnu}$: 1.97 $\mathsf{react}$: 2.0 $\mathsf{react}$: 2.0 $\mathsf{react}$: 1.94
[150] -0.264 (0.792) [138] -514.302 (*) [148] -547.413 (*) [160] -14437.976 (*) [152] -14871.238 (*) [140] -397.488 (*)
$\mathsf{pro}_{\mathsf{SAA}}$: 0.99 $\mathsf{pro}_{0.9}$: 0.02 $\mathsf{pro}_{\mathsf{SAA}}$: 0.02 $\mathsf{pro}_{0.9}$: 0.0 $\mathsf{pro}_{\mathsf{SAA}}$: 0.0 $\mathsf{stnu}$: 0.07
$\mathsf{pro}_{0.9}$: 1.01 $\mathsf{stnu}$: 1.98 $\mathsf{stnu}$: 1.98 $\mathsf{react}$: 2.0 $\mathsf{react}$: 2.0 $\mathsf{react}$: 1.93
[150] -0.707 (0.481) [148] 2600.902 (*) [146] 2585.661 (*) [155] -25795.773 (*) [145] -23210.69 (*) [141] -121.392 (*)
$\mathsf{pro}_{0.9}$: 0.96 $\mathsf{stnu}$: 1.99 $\mathsf{stnu}$: 1.99 $\mathsf{pro}_{0.9}$: 0.0 $\mathsf{pro}_{\mathsf{SAA}}$: 0.0 $\mathsf{stnu}$: 0.14
$\mathsf{pro}_{\mathsf{SAA}}$: 1.04 $\mathsf{pro}_{0.9}$: 0.01 $\mathsf{pro}_{\mathsf{SAA}}$: 0.01 $\mathsf{react}$: 2.0 $\mathsf{react}$: 2.0 $\mathsf{react}$: 1.86
[94] 0.141 (0.888) [114] 21743.706 (*) [114] 21911.639 (*) [99] -46429.42 (*) [76] -40238.643 (*) [94] -41.405 (*)
$\mathsf{pro}_{\mathsf{SAA}}$: 1.0 $\mathsf{stnu}$: 2.0 $\mathsf{stnu}$: 2.0 $\mathsf{pro}_{0.9}$: 0.0 $\mathsf{pro}_{\mathsf{SAA}}$: 0.0 $\mathsf{stnu}$: 0.21
$\mathsf{pro}_{0.9}$: 1.0 $\mathsf{pro}_{0.9}$: 0.0 $\mathsf{pro}_{\mathsf{SAA}}$: 0.0 $\mathsf{react}$: 2.0 $\mathsf{react}$: 2.0 $\mathsf{react}$: 1.79
Table 23: Magnitude t-test on time online for noise factor c=2 (double hits). Each cell shows on the first row [nr pairs] t-stat (p-value) with (*) for p $<$ 0.05 and on the second row the normalized average of method 1 and on the third row the normalized average of method 2.

### v2

@@ Appendix F Results Tables Partial Orderings
Wilcoxon: [n] z (p) [288] -8.783 (*) [285] -10.58 (*) [284] -13.068 (*) [122] -4.786 (*) [146] -5.008 (*) [141] -1.441 (0.15)
proportion: [n] z (p) [288] 0.576 (*) [285] 0.653 (*) [284] 0.82 (*) [23] 1.0 (*) [146] 0.836 (*) [141] 0.702 (*)
Table 23: Pairwise comparison on time offline for noise factor $\epsilon=2$. Using a Wilcoxon test and a proportion test. Including all instances for which at least one of the two methods found a feasible solution. Note that the ordering matters: the first method showed is the better of the two in each pair method 1 - method 2 according to Wilcoxon. Each cell shows on the first row [nr pairs] z-value (p-value) of the Wilcoxon test with (*) for p $<$ 0.05. Each cell shows on the second row [nr pairs] ratio of wins (p-value) with (*) for p $<$ 0.05
@@
Wilcoxon: [n] z (p) [400] -0.719 (0.472) [400] -17.331 (*) [400] -17.331 (*) [400] -17.331 (*) [400] -17.331 (*) [400] -12.488 (*)
proportion: [n] z (p) [391] 0.453 (0.069) [400] 1.0 (*) [400] 1.0 (*) [400] 1.0 (*) [400] 1.0 (*) [400] 0.925 (*)
Table 24: Pairwise comparison on time online for noise factor $\epsilon=1$. Using a Wilcoxon test and a proportion test. Including all instances for which at least one of the two methods found a feasible solution. Note that the ordering matters: the first method showed is the better of the two in each pair method 1 - method 2. Each cell shows on the first row [nr pairs] z-value (p-value) of the Wilcoxon test with (*) for p $<$ 0.05. Each cell shows on the second row [nr pairs] proportion (p-value) with (*) for p $<$ 0.05
@@
Wilcoxon: [n] z (p) [235] -0.384 (0.701) [258] -9.554 (*) [258] -9.212 (*) [234] -13.262 (*) [235] -12.623 (*) [258] -8.801 (*)
proportion: [n] z (p) [66] 0.47 (0.712) [258] 0.907 (*) [258] 0.899 (*) [234] 1.0 (*) [235] 0.987 (*) [258] 0.891 (*)
@@
Wilcoxon: [n] z (p) [182] -1.204 (0.229) [216] -4.456 (*) [217] -4.913 (*) [177] -11.279 (*) [182] -10.957 (*) [217] -11.457 (*)
proportion: [n] z (p) [79] 0.43 (0.261) [216] 0.815 (*) [217] 0.825 (*) [177] 0.994 (*) [182] 0.984 (*) [217] 0.968 (*)
@@
Wilcoxon: [n] z (p) [146] -0.482 (0.63) [288] -8.467 (*) [284] -9.06 (*) [122] -9.585 (*) [141] -6.136 (*) [285] -14.285 (*)
proportion: [n] z (p) [144] 0.458 (0.359) [288] 0.576 (*) [284] 0.585 (*) [122] 1.0 (*) [141] 0.837 (*) [285] 0.982 (*)
Table 25: Pairwise comparison on time online for noise factor $\epsilon=2$. Using a Wilcoxon test and a proportion test. Including all instances for which at least one of the two methods found a feasible solution. Note that the ordering matters: the first method showed is the better of the two in each paier method 1 - method 2. Each cell shows on the first row [nr pairs] z-value (p-value) of the Wilcoxon test with (*) for p $<$ 0.05. Each cell shows on the second row [nr pairs] proportion (p-value) with (*) for p $<$ 0.05